A bag contains 5 lemon gumdrops and 7 cherry gumdrops. Four gumdrops are selected at random.What is the probability that two of each flavor will be selected? Round your answer to the nearest thousandth.
Answer by stanbon(75887)

A bag contains 5 lemon gumdrops and 7 cherry gumdrops. Four gumdrops are selected at random.What is the probability that two of each flavor will be selected? Round your answer to the nearest thousandth.
------------
# of ways to pick 2 lemon: 5C2 = 10
# of ways to pick 2 cherry:7C2 = 21
---------------------------
# of ways to pick 4 randomly: 12C4 = (12*11*10*8)/(1*2*3*4) = 495
------------
P(2 lemon and 2 cherry) = [10*21]/495 = 0.4242
======================
Cheers,
Stan H.